Two e-jave lengths of >,, 2 ave sevt hrough a Youms dauble -slit dlevice smul tone ou is to coincide cu lth He 2 order 2
Answer #1

a)
use,
Y = (D*n*(wavelength))/d

so, for wavelength = lameda1
n = 3
Y1 = (D*3*(lameda1))/d

similarly,
for wavelngth = lameda2
n=4
Y2 = (D*4*(lameda2))/d

It is given that
Y1 = Y2
(D*3*(lameda1))/d = D*4*(lameda2))/d
lameda1 = (4/3)*(lameda2)

b)
given,
lameda1 = 600 nm
so,
600 = = (4/3)*(lameda2)
lameda2 = (3/4)*600
= 450 nm